OpenStudy (anonymous):

Let f(x) = x2 - 16. Find f-1(x).

OpenStudy (anonymous):

let y be f(x)

OpenStudy (anonymous):

\[ x=f(f^{-1}(x))=[f^{-1}(x)]^2-16 \]

OpenStudy (anonymous):

therefor y=x2-16

OpenStudy (anonymous):

y+16=x2

OpenStudy (anonymous):

@Nurialozza96 what would be the first step?

OpenStudy (anonymous):

root(y+16)=y therefore f-1(x)=root(x-16)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@UditKulka Are you trying to prevent learning?

OpenStudy (anonymous):

@Nurialozza96 Do you still need help?

OpenStudy (anonymous):

A little, Im not so sure of what im doing here /:

OpenStudy (anonymous):

\[ x=[f^{-1}(x)]^2-16 \] First... do you want to add multiply or what?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

There is nothing to multiply right now.

OpenStudy (anonymous):

Actually when isolating, the first thing you want to do is add/subtract.

OpenStudy (anonymous):

+16?

OpenStudy (anonymous):

\[ x+16=[f^{-1}(x)]^2 \]

OpenStudy (anonymous):

thren im guessing you square root it

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I got it ! is there a rule behind these problems ? seems to be ..

OpenStudy (anonymous):

Okay do you know the order of operations?

OpenStudy (anonymous):

Paranthesis exponent multiply divide add subtract ?

OpenStudy (anonymous):

Yes. When doing these problems, you want to do it in reverse order.

OpenStudy (anonymous):

ohh ok ok (: thankyou

OpenStudy (anonymous):

what about Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)

OpenStudy (anonymous):

@henryarias5 Ask a separate question.

OpenStudy (anonymous):

Anyway, in this case we have: \[ |f^{-1}(x)|=\sqrt{x+16} \]

OpenStudy (anonymous):

This means that: \[ f^{-1}(x)=\sqrt{x+16}\quad\text{and} \quad -\sqrt{x+16} \]

OpenStudy (anonymous):

This is NOT a function because it has two outputs. Technically this function does not have an inverse function.